OpenStudy (anonymous):

If sin Θ = 2 over 3 and tan Θ < 0, what is the value of cos Θ?

OpenStudy (anonymous):

\[\sin \theta=\frac{ 2 }{ 3 },\tan \theta <0\] hence theta lies in second quadrant ,where cos theta is negative.

OpenStudy (anonymous):

\[\cos \theta=-\sqrt{1-\sin ^2\theta }\]

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

\[\cos ^2 \theta +\sin ^2 \theta=1\] \[\cos ^ 2 \theta =1-\sin ^2 \theta \] \[\cos \theta =\pm \sqrt{1-\sin ^2 \theta }\] as \[\sin \theta =\frac{ 2 }{ 3 }>0,\tan \theta <0 \] \[\sin \theta >0 shows ~that~\theta~lies~\in~1st~or~2nd ~quadrant.\] \[\tan \theta <0,\theta \in ~2nd~or~4th~quadrant.\] combining theta lies in 2nd quadrant. \[\cos \theta=-\sqrt{1-\sin ^2 \theta }=-\sqrt{1-\left( \frac{ 2 }{ 3 } \right)^2}=?\]